OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? 2x-y=-5 x+3y=22 A. (1, 7) B. (4, 9) C. (7, 19) D. (8, 21) @ganeshie8 @hartnn @mathstudent55 @Abhisar @dan815 @iambatman @Callisto @.Sam.

OpenStudy (mathstudent55):

Multiply the first equation by 3 and add tot eh second equation. Solve for x. Then find the answer that has that same x value.

OpenStudy (anonymous):

2x-y=-5 -----(1) x+3y=22 -----(2) Multiply second eq with 2 2(x+3y=22) 2x +6y = 44 ------(3) subtract (1) from (3) (2x +6y) - (2x -y) = 44- (-5) 7y = 49 find x and y

OpenStudy (anonymous):

is it a @sangya21 @mathstudent55

OpenStudy (anonymous):

Correct. :)

OpenStudy (mathstudent55):

2x-y=-5 --- *3 ---> 6x - 3y = -15 x+3y=22 ---------> x + 3y = 22 ----------- 7x = 7 x = 1 Since the only choice that has x = 1 is A, the answer must be A, and don't even need to solve for y.

OpenStudy (mathstudent55):

You are correct.

OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? x+3=12 -4x+3y=-3 A. (0, 4) B. (6, 2) C. (3, 3) D. (0, −1)

OpenStudy (anonymous):

Subtract 2 from 1. Find the value of x and y

OpenStudy (anonymous):

is this d??

OpenStudy (anonymous):

Correct :)

OpenStudy (anonymous):

omg are u serious thanks

OpenStudy (anonymous):

Anytime :)

OpenStudy (anonymous):

can u help me with 2 more?

OpenStudy (anonymous):

sure